Suppose that 3% of computer chips produced by a certain machine are defective. The chips are put into packages of 20 chips for distribution to retailers. What is the probability that a randomly selected package of chips will contain at least 2 defective chips?
1
Expert's answer
2020-05-15T15:10:23-0400

n=20,p=0.03

P(x≥2)=1-P(x=0)-P(x=1)

P(x=0)=C20,0*0.030*(1-0.03)20-0=0.5438

P(x=1)=C20,1*0.031*0.9719=0.3363

P(x≥2)=1-0.5438-0.3363=0.12